# SDEs with critical general distributional drifts: sharp solvability and blow-ups

### v4

@@ Abstract.
We establish weak well-posedness for SDEs having discontinuous diffusion coefficients and general distributional drifts that may introduce blow-up effects. Our drifts satisfy minimal assumptions, i.e. we assume only that the Cauchy problem for the Kolmogorov backward equation is well-posed in the standard Hilbert triple $W^{1,2}\hookrightarrow L^{2}\hookrightarrow W^{-1,2}$. By a result of Mazya and Verbitsky, these assumptions are precisely those drifts that can be represented as the sum of a form-bounded component (encompassing, for example, Morrey or Chang-Wilson-Wolff drifts) and a divergence-free distributional component in the ${\rm BMO}^{-1}$ space of Koch and Tataru.
@@
Stochastic differential equations, singular and distributional drifts, divergence-free drifts, form-boundedness, ${\rm BMO}^{-1}$, De Giorgi’s method, particle systems, Keller-Segel model, blow-ups
@@ 1. Introduction
with critical general distributional drift $c$ and discontinuous diffusion coefficients $\sigma$ (see Section 5 and Section 7 for our precise setting). Here $\{B_{t}\}_{t\geq 0}$ denotes a $d$-dimensional Brownian motion. SDEs with singular drifts arise in various physical models, for instance, the passive-tracer model, where a time-dependent drift $c$ is the velocity field obtained from the Navier–Stokes equations [MK], or finite-particle approximations of the Keller–Segel model of chemotaxis [CP, FJ]. In the latter case one observes blow-ups, i.e. when even the weak existence for SDE (1.1) fails once one replaces $c$ by $(1+\varepsilon)c$, $\varepsilon>0$, in which case all particles collide a.s. in finite time and can stay “stuck” indefinitely. Both of these models are within the scope of the present work.
We postpone a survey of the recent literature on SDEs with singular drifts (much of it produced in the past decade) until after we have stated and discussed our assumptions on $c$.
@@ 1.2. Result #2
where the arrow $\rightarrow$ in this diagram means inclusion. (These inclusions were proved in [M, Theorem 1.4.7] and [MV2, Theorem V]. See discussion before Theorem 6.3 and the proof in Appendix D.) Choosing the inclusion into $\mathbf{MF}_{\delta}$ leads to an approach to studying the Kolmogorov backward equation based on “Caccioppoli’s iterations” for establishing the classical Caccioppoli’s inequality [KV], see Remark 6.1 for details; choosing the inclusion into $\mathbf{BMO}^{-1}$ amounts to absorbing the stream matrix $Q$ (for $b=\nabla Q$) into the diffusion coefficients and obtaining a Caccioppoli-type inequality from there [H, SSŠZ]. We discuss this in Section 6.
@@ 1.4. Literature
2) If we restrict attention to Borel-measurable drifts, the most closely related results are the recent works of Krylov [Kr1, Kr2, Kr3, Kr4, Kr5] and of Röckner and Zhao [RZ]. In [Kr1], Krylov treats a class of diffusion coefficients much larger than ours (his diffusion coefficients are in $\rm VMO$} class, or have small $\rm BMO$} norm), but a smaller class of drifts (namely, the Morrey class $M_{(\frac{d}{2}\vee 2)+\varepsilon}$). The irregular diffusion coefficients require estimates on second-order derivatives of solutions to the Kolmogorov backward equation, whereas we neither assume nor have such estimates under our assumptions on the drift. Krylov’s drifts can also produce blow-up phenomena, but he assumes that the size of the singularity, measured by the Morrey norm, remains below a small, dimension-dependent constant. By contrast, our goal is to reach the maximal admissible strength of singularity (in Example 1.1, the interaction strength $\kappa$; or the supremum (1.2)). Other recent results of Krylov are described at various points in this work.
@@ 5. General drifts
    (A relaxed approximation uniqueness) The limit in (i) does not depend on the choice of $\{b_{n}\}\in[b]$ and $\{q_{m}\}\in[q]$. Furtermore, if $\delta<1$ and $\{b_{n}\}\in[b]\cap[L^{2}]^{d}$, then already weak convergence
@@
    Thus, when $\delta<1$, we can extend quite substantially the class of admissible approximations of $b$.
@@
    $X_{t}=x-\int_{0}^{t}b(X_{s})ds+\sqrt{2}B_{t}.$ (5.4)
    That is, $X=\Phi(B)$ for some Borel measurable function $\Phi$.
@@ 8.0.4. Proof of $3^{\circ}$)
$=:I_{1}+I_{2}$ (8.21)
By Step 1 and $\rho\leq\sqrt{\rho}$, term $I_{1}$ can be estimated as follows: for every $\varepsilon>0$, for all $R=R(\varepsilon)>0$ sufficiently large
@@ 11.2. Global weighted $L^{2}$ summability of a form-bounded drift
We have $k_{0}=0$, but we will not encounter the volume of $\{u>0\}$ in the proof (clearly, $|\{u>0\}|$ can be infinite).
@@
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{p\theta}^{p\theta}\leq\tilde{C}\frac{1}{\xi^{p\theta}}(\mu-\mu_{0})^{-\beta\theta}H^{\theta}|\{u>k_{m+1}\}|^{2-\frac{\theta}{\theta_{0}}}.$
@@
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{L^{p\theta}(B^{m+1})}^{p\theta}\leq\tilde{C}|\{u>k_{m+1}\}|^{2-\frac{\theta}{\theta_{0}}}.$
@@ Appendix D Multiplicative form-boundedness and Morrey class $M_{1}$
The following result was proved in [M, Theorem 1.4.7]. We reproduce the proof for reader’s convenience.
@@
for some constant $K$ independent of $r>0$ and $x\in\mathbb{R}^{d}$ (then $K$ is proportional to $\delta$).
In other words, $b\in\mathbf{MF}_{\delta}$ with $c_{\delta}=0$ if and only if $|b|$ belongs to the Morrey class $M_{1}$.
Proof of Proposition D.1.
Define weighted Lebesgue norm over ball $B_{r}(x)$:
$\|\varphi\|^{q}_{L^{q}(B_{r}(x),|b|)}:=\int_{B_{r}(x)}|\varphi(y)|^{q}|b(y)|dy.$
Define analogously $L^{q}(\mathbb{R}^{d},|b|)$. D.R. Adams’ Lemma C.2 will play a crucial role in the proof.
Proof of (D.2) $\Rightarrow$ (D.1). So, we assume that (D.2) holds, and prove the multiplicative inequality (D.1).
Lemma D.1.
For every $\varphi\in C(\overline{B_{r}})$,
$\|\varphi\|_{L^{2}(B_{r},|b|)}\leq C_{2}\left(r^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r})}+r^{-\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r})}\right),$ (D.3)
for constant $C_{2}$ independent of $\varphi$ or $r$.
@@
Any function $\varphi\in C^{\infty}(\overline{B_{1}})$ can be extended to a function in $C_{c}^{1}(B_{2})$, which we still denote, with some abuse of notation, by $\varphi$, so that
$\|\nabla\varphi\|_{L^{2}(B_{2})}\leq C_{1}\bigl(\varphi\|_{L^{2}(B_{1})}+\|\nabla\varphi\|_{L^{2}(B_{1})}\bigr),$
where $C_{1}$ does not depend on $\varphi$. After rescaling, we obtain
$\|\nabla\varphi\|_{L^{2}(B_{2r})}\leq C\left(\|\nabla\varphi\|_{L^{2}(B_{r})}+r^{-1}\|\varphi\|_{L^{2}(B_{r})}\right).$ (D.4)
Now that the integrals over $B_{r}$ do not depend on the choice of the extension of $\varphi$. The integrals over $B_{2r}$which do depend on the choice of the extension, like the one in the left-hand side of the previous inequality, will appear below, but only at the intermediate steps, and will not appear in the final estimates.
Now, put $q=2\frac{d-1}{d-2}$. Then our hypothesis $b\in M_{1}$ allows us to apply Lemma C.2 with thus selected $q$ and $p=2$. For $p=2$, we have $\|(-\Delta)^{\frac{1}{2}}u\|_{2}=\|\nabla u\|_{2}$. So,
$\biggl(\|\varphi\|_{L^{q}(B_{r},|b|)}\leq\biggr)\quad\|\varphi\|_{L^{q}(B_{2r},|b|)}\leq C\|\nabla\varphi\|_{L^{2}(B_{2r})},$ (D.5)
where, of course, it is crucial that $\varphi$ has support in the open ball $B_{2r}$, so integrating over $\mathbb{R}^{d}$ is equivalent to integrating over $B_{2r}$.
By Hölder’s inequality,
$\|\varphi\|_{L^{2}(B_{r},|b|)}\leq\langle|b|\mathbf{1}_{B_{r}}\rangle^{\frac{1}{2}-\frac{1}{q}}\|\varphi\|_{L^{q}(B_{r},|b|)},$ (D.6)
where $\frac{1}{2}-\frac{1}{q}=\frac{1}{2}\frac{1}{d-1}$.
Finally, assemblying estimates (D.6)-(D.4) and using our hypothesis (D.2) on $|b|$, we arrive at
$\|\varphi\|_{L^{2}(B_{r},|b|)}$ $\leq\langle|b|\mathbf{1}_{B_{r}}\rangle^{\frac{1}{2}-\frac{1}{q}}\|\varphi\|_{L^{q}(B_{r},|b|)}\leq CK^{\frac{1}{2}\frac{1}{d-1}}r^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{2r})}$
$\leq C_{2}\left(r^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r})}+r^{-\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r})}\right),$
i.e. we have proved (D.3). ∎
We are in position to complete the proof of implication (D.2) $\Rightarrow$ (D.1). Fix some $r_{0}>0$, and do the following for every $x\in\mathbb{R}^{d}$:
– If $r_{0}\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}\geq\|\varphi\|_{L^{2}(B_{r_{0}}(x))}$, then from (D.3)
$\|\varphi\|_{L^{2}(B_{r_{0}}(x),|b|)}$ $\leq C_{2}\left(r_{0}^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}+\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r_{0}}(x))}^{\frac{1}{2}}\right)$
$\leq C_{2}\left(r_{0}^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}+\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r_{0}}(x))}^{\frac{1}{2}}\right).$ (D.7)
- Otherwise, increase $r$ until one reaches equality $r\|\nabla\varphi\|_{L^{2}(B_{r}(x))}=\|\varphi\|_{L^{2}(B_{r}(x))}$, in which case (D.3) yields
$\|\varphi\|_{L^{2}(B_{r}(x),|b|)}$ $\leq C_{2}\left(\|\nabla\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}+\|\nabla\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\right)$
$=2C_{2}\|\nabla\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}.$ (D.8)
In any case, we obtain, for every $x\in\mathbb{R}^{d}$ and $r=r(x)$ (either $r_{0}$ or increased $r$, depending on $x$):
$\|\varphi\|_{L^{2}(B_{r}(x),|b|)}\leq 2C_{2}\bigg(\|\nabla\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}+r_{0}^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}\bigg).$ (D.9)
Our goal now is to turn local estimate (D.9) into the multiplicative inequality (D.1). So, we cover $\mathrm{supp}|b|$ with balls $\{B_{r(x)}(x)\}$. Using [M, Theorem 1.2.1], i.e. a variant of the Besicovich covering theorem, one can extract at most countable subcover of finite multiplicity $M=M(d)$. We emphasize that while the covering depends on $\varphi$, the multiplicity $M$ does not. Let $\{B^{i}\}$ be this sub-cover. Squaring (D.9) (considered over each ball $B^{i}$ in the sub-cover), summing up in $i$, and using Hölder inequality, we have:
$\|\varphi\|^{2}_{L^{2}(\mathbb{R}^{d},|b|)}$ $\leq M\sum_{i}\|\varphi\|^{2}_{L^{2}(B^{i},|b|)}$
@@
[…]
By Jensen’s inequality,
$\big(\sum_{i}\|\nabla\varphi\|_{L^{2}(B^{i})}^{2}\big)^{1/2}\leq\sum_{i}\|\nabla\varphi\|_{L^{2}(B^{i})}\leq M\|\nabla\varphi\|_{2}$
and similarly for $\|\varphi\|_{2}$, which gives us
$\|\varphi\|^{2}_{L^{2}(\mathbb{R}^{d},|b|)}\leq M^{2}C_{3}(\|\nabla\varphi\|_{2}\|\varphi\|_{2}+r_{0}\|\nabla\varphi\|^{2}_{2})$
Since $r_{0}$ was fixed arbitrarily, letting $r_{0}\downarrow 0$ yields the multiplicative inequality (D.1).
The reverse direction (D.1) $\Rightarrow$ (D.2) is easier. Choosing in (D.1) test functions $\varphi=\varphi_{r}\in C_{c}^{\infty}(B_{r}(x))$ such that $\|\nabla\varphi\|_{2}\leq cr^{\frac{d-2}{2}}$, $\|\varphi\|_{2}\leq cr^{d/2}$, we obtain $\langle|b|,\mathbf{1}_{B_{r}(x)}\rangle\leq Cr^{d-1}$, i.e. $b\in M_{1}$. This completes the proof. ∎

### v5

@@ Abstract.
We establish weak well-posedness for SDEs having discontinuous diffusion coefficients and general distributional drifts that may introduce local blow-up effects. Our drifts satisfy minimal assumptions, i.e. we assume only that the Cauchy problem for the Kolmogorov backward equation is well-posed in the standard Hilbert triple $W^{1,2}\hookrightarrow L^{2}\hookrightarrow W^{-1,2}$. By a result of Mazya and Verbitsky, these assumptions are precisely those drifts that can be represented as the sum of a form-bounded component (encompassing, for example, Morrey or Chang-Wilson-Wolff drifts) and a divergence-free distributional component in the ${\rm BMO}^{-1}$ space of Koch and Tataru.
@@
Stochastic differential equations, singular and distributional drifts, divergence-free drifts, form-boundedness, ${\rm BMO}^{-1}$, De Giorgi’s method, particle systems, Keller-Segel model, local blow-ups
@@ 1. Introduction
with critical general distributional drift $c$ and discontinuous diffusion coefficients $\sigma$ (see Section 5 and Section 7 for our precise setting). Here $\{B_{t}\}_{t\geq 0}$ denotes a $d$-dimensional Brownian motion. SDEs with singular drifts arise in various physical models, for instance, the passive-tracer model, where a time-dependent drift $c$ is the velocity field obtained from the Navier–Stokes equations [MK], or finite-particle approximations of the Keller–Segel model of chemotaxis [CP, FJ]. In the latter case one observes local blow-ups, i.e. when even the weak existence for SDE (1.1) fails once one replaces $c$ by $(1+\varepsilon)c$, $\varepsilon>0$, in which case all particles collide a.s. in finite time and can stay “stuck” indefinitely. Both of these models are within the scope of the present work.
We postpone a survey of the recent literature on SDEs with singular drifts until after we have stated and discussed our assumptions on $c$.
@@ 1.2. Result #2
where the arrow $\rightarrow$ in this diagram means inclusion. (These inclusions were proved in [M, Theorem 1.4.7] and [MV2, Theorem V]. See discussion before Theorem 6.3 and Krylov’s proof in Appendix D.) Choosing the inclusion into $\mathbf{MF}_{\delta}$ leads to an approach to studying the Kolmogorov backward equation based on “Caccioppoli’s iterations” for establishing the classical Caccioppoli’s inequality [KV], see Remark 6.1 for details; choosing the inclusion into $\mathbf{BMO}^{-1}$ amounts to absorbing the stream matrix $Q$ (for $b=\nabla Q$) into the diffusion coefficients and obtaining a Caccioppoli-type inequality from there [H, SSŠZ]. We discuss this in Section 6.
@@ 1.4. Literature
2) If we restrict attention to Borel-measurable drifts, the most closely related results are the recent works of Krylov [Kr1, Kr2, Kr3, Kr4, Kr5] and of Röckner and Zhao [RZ]. In [Kr1], Krylov treats a class of diffusion coefficients much larger than ours (his diffusion coefficients are in $\rm VMO$} class, or have small $\rm BMO$} norm), but a smaller class of drifts (namely, the Morrey class $M_{(\frac{d}{2}\vee 2)+\varepsilon}$). The irregular diffusion coefficients require estimates on second-order derivatives of solutions to the Kolmogorov backward equation, whereas we neither assume nor have such estimates under our assumptions on the drift. Krylov’s drifts can also produce blow-up phenomena, but he assumes that the size of the singularity, measured by the Morrey norm, remains below a small, dimension-dependent constant. By contrast, our goal is to reach the maximal admissible strength of singularity (in Example 1.1, the interaction strength $\kappa$; or the supremum (1.2)). Other recent results of Krylov are described below.
@@ 5. General drifts
    (A relaxed approximation uniqueness) The limit in (i) does not depend on the choice of $\{b_{n}\}\in[b]$ and $\{q_{m}\}\in[q]$. Furtermore, if $\delta<1$ and $\{b_{n}\}\in[b]\cap[L^{2}]^{d}$, then already the weak convergence
@@
    Thus, when $\delta<1$, we can extend the class of admissible approximations of $b$.
@@
    $X_{t}=x-\int_{0}^{t}b(X_{s})ds+\sqrt{2}B_{t}$ (5.4)
    (so, $X=\Phi(B)$ for a Borel measurable function $\Phi$; the solution depends only on the Brownian motion).
@@ 8.0.4. Proof of $3^{\circ}$)
$=:I_{1}+I_{2}$
By Step 1 and $\rho\leq\sqrt{\rho}$, the term $I_{1}$ can be estimated as follows: for every $\varepsilon>0$, for all $R=R(\varepsilon)>0$ sufficiently large
@@ 11.2. Global weighted $L^{2}$ summability of a form-bounded drift
We have $k_{0}=0$, but we will not encounter the volume of $\{w>0\}$ in the proof (clearly, $|\{w>0\}|$ can be infinite).
@@
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{p\theta}^{p\theta}\leq\tilde{C}\frac{1}{\xi^{p\theta}}(\mu-\mu_{0})^{-\beta\theta}H^{\theta}|\{w>k_{m+1}\}|^{2-\frac{\theta}{\theta_{0}}}.$
@@
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{L^{p\theta}(B^{m+1})}^{p\theta}\leq\tilde{C}|\{w>k_{m+1}\}|^{2-\frac{\theta}{\theta_{0}}}.$
@@ Appendix D Multiplicative form-boundedness and Morrey class $M_{1}$
for some constant $K$ independent of $r>0$ and $x\in\mathbb{R}^{d}$, i.e. $|b|\in M_{1}$; then $K$ is proportional to $\delta$.
This result was proved by Mazya [M, Theorem 1.4.7]. The proof in [M] uses the Besicovich covering theorem. Below we give a shorter and elementary proof due to Krylov [Kr7].
@@
Let us prove (D.2) $\Rightarrow$ (D.1).
Step 1.  By Adams’ Lemma C.2,
$\langle|b|,|\varphi|^{q}\rangle^{\frac{1}{q}}\leq CK\|\nabla\varphi\|_{2},\quad q=2\frac{d-1}{d-2}>2,$
for every $\varphi\in C_{c}^{\infty}$. Hence, by Hölder’s inequality, for $\varphi$ with support in a ball of radius $1$,
$\langle|b|,\varphi^{2}\rangle$ $\leq\langle|b|\mathbf{1}_{B_{1}}\rangle^{\frac{q-2}{q}}\langle|b|,|\varphi|^{q}\rangle^{\frac{2}{q}}$
$\leq C_{0}K\|\nabla\varphi\|^{2}_{2}.$
Using the cutoff function $\zeta_{x}$ with support in the ball of radius $1$ centered at some $x$, we now obtain, for arbitrary $\varphi\in C_{c}^{\infty}$,
$\langle|b|,\varphi^{2}\zeta_{x}^{2}\rangle\leq C_{1}K\|(\nabla\varphi)\zeta_{x}\|^{2}_{2}+C_{1}K\|(\nabla\zeta_{x})\varphi\|_{2}^{2}.$
We now sum up over all $x\in\mathbb{Z}^{d}$, obtaining
$\langle|b|,\varphi^{2}\rangle\leq C_{2}K\big(\|\nabla\varphi\|^{2}_{2}+\|\varphi\|_{2}^{2}\big).$ (D.3)
Step 2. Put $b_{\lambda}(x):=b(x/\lambda)$, $\lambda>0$. The Morrey norm $\|\cdot\|_{M_{1}}$ scales as follows:
$\|b_{\lambda}\|_{M_{1}}=\lambda\|b\|_{M_{1}},$
so
$\langle|b_{\lambda}|\mathbf{1}_{B_{r}(x)}\rangle\leq K\lambda r^{d-1},\quad x\in\mathbb{R}^{d},\;r>0.$
Therefore, by (D.3),
$\langle|b_{\lambda}|,\varphi^{2}\rangle\leq C_{2}K\lambda\big(\|\nabla\varphi\|^{2}_{2}+\|\varphi\|_{2}^{2}\big),\quad\lambda>0.$ (D.4)
We now deduce the sought inequality (D.1) from the family of inequalities (D.4). Put $\psi(\cdot):=\varphi(\lambda\cdot)$. Then
$\langle|b|,\psi^{2}\rangle$ $=\lambda^{-d}\langle|b_{\lambda}|,\varphi^{2}\rangle$
@@
$\leq\lambda^{-d}C_{2}K\lambda\big(\|\nabla\varphi\|^{2}_{2}+\|\varphi\|_{2}^{2}\big)$
$=C_{2}K\big(\lambda^{-1}\|\nabla\psi\|^{2}_{2}+\lambda\|\varphi\|_{2}^{2}\big).$
It remains to minimize the right-hand side in $\lambda$, i.e. take $\lambda=\frac{\|\nabla\psi\|_{2}}{\|\psi\|_{2}},$ to obtain (D.1) with $\delta=2C_{2}K$.
The reverse direction (D.1) $\Rightarrow$ (D.2) is easier. Choosing in (D.1) test functions $\varphi=\varphi_{r}\in C_{c}^{\infty}(B_{r}(x))$ such that $\|\nabla\varphi\|_{2}\leq cr^{\frac{d-2}{2}}$, $\|\varphi\|_{2}\leq cr^{d/2}$, we obtain $\langle|b|,\mathbf{1}_{B_{r}(x)}\rangle\leq Cr^{d-1}$, i.e. $|b|\in M_{1}$. This completes the proof. ∎